# Joint Cooperative and Non-Cooperative Localization in WSNs with Distributed Scaled Proximal ADMM Algorithms

### v1

@@ VI Conclusion
Appendix A Solution to Remark 1
@@
By substituting the explicit form of equation (III) into problem (31), the optimization problem (31) can be reformulated as:
$\min_{\mathbf{z}\in\mathcal{X}}\sum_{i\in\mathcal{N}}\Big($ $[(c+1)N_{i}+1]\|\mathbf{x}_{i}-\tilde{\mathbf{x}}_{i}^{t+1}\|^{2}+\sum_{j\in\mathcal{N}_{i}}c\|\mathbf{p}_{i,j}^{-}-(\tilde{\mathbf{p}}_{i,j}^{-})^{t+1}\|^{2}$
$+\sum_{j\in\mathcal{N}_{i}}\|\mathbf{p}_{i,j}^{+}-(\tilde{\mathbf{p}}_{i,j}^{+})^{t+1}\|^{2}+(2cN_{i}+1)\|\mathbf{y}_{i}-\tilde{\mathbf{y}}_{i}^{t+1}\|^{2}$
$+\sum_{j\in\mathcal{N}_{i}}c\|\mathbf{q}_{i,j}^{-}-(\tilde{\mathbf{q}}_{i,j}^{-})^{t+1}\|^{2}+\sum_{j\in\mathcal{N}_{i}}c\|\mathbf{q}_{i,j}^{+}-(\tilde{\mathbf{q}}_{i,j}^{+})^{t+1}\|^{2}\Big).$
@@
The optimal solution for $\mathbf{y}_{i}^{t+1}$ is:
$\mathbf{y}_{i}^{t+1}=\tilde{\mathbf{y}}_{i}^{t+1},\text{for }i\in\mathcal{N}$
@@
The terms related to $\mathbf{q}_{i,j}^{-}$ are:
$c\|\mathbf{q}_{i,j}^{-}-(\tilde{\mathbf{q}}_{i,j}^{-})^{t+1}\|^{2}+c\|\mathbf{q}_{i,j}^{-}-(\tilde{\mathbf{q}}_{j,i}^{+})^{t+1}\|^{2},\quad\forall i\in\mathcal{N},j\in\mathcal{N}_{i}.$
The terms related to $\mathbf{q}_{i,j}^{+}$ are:
$c\|\mathbf{q}_{i,j}^{+}-(\tilde{\mathbf{q}}_{i,j}^{-})^{t+1}\|^{2}+c\|\mathbf{q}_{i,j}^{+}-(\tilde{\mathbf{q}}_{j,i}^{+})^{t+1}\|^{2},\quad\forall i\in\mathcal{N},j\in\mathcal{N}_{i}.$
@@
For $i\in\mathcal{N},j\in\mathcal{N}$, the closed-form solution for $\mathbf{q}_{i,j}^{+}$ is:
$(\mathbf{q}_{i,j}^{-})^{t+1}=\frac{1}{2}\left((\tilde{\mathbf{q}}_{i,j}^{-})^{t+1}+(\tilde{\mathbf{q}}_{j,i}^{+})^{t+1}\right),\quad\forall i\in\mathcal{N},j\in\mathcal{N}_{i}.$
For $i\in\mathcal{N},j\in\mathcal{N}$, the closed-form solution for $\mathbf{q}_{i,j}^{-}$ is:
$(\mathbf{q}_{i,j}^{+})^{t+1}=\frac{1}{2}\left((\tilde{\mathbf{q}}_{i,j}^{+})^{t+1}+(\tilde{\mathbf{q}}_{j,i}^{-})^{t+1}\right),\quad\forall i\in\mathcal{N},j\in\mathcal{N}_{i}.$
Appendix B (Proof of Lemma1)
@@
$\quad+\frac{3\left(N_{\max}+2\right)}{c}\|\mathbf{H}_{i}\Delta\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{D}_{i}\Delta\mathbf{w}_{i}^{t}\|^{2}$
$\quad+3c\left(2N_{\max}+1\right)\|\Delta\tilde{\mathbf{z}}_{i}^{t+1}-\Delta\mathbf{z}_{i}^{t+1}\|_{\mathbf{A}_{i}^{T}\mathbf{A}_{i}}^{2}$
$\quad+3\left(1+2c\right)\left(2+N_{\max}\right)\|\Delta\tilde{\mathbf{z}}_{i}^{t+1}-\Delta\mathbf{z}_{i}^{t}\|_{\mathbf{B}_{i}^{T}\mathbf{B}_{i}}^{2}\Bigg],$ (49)
where $N_{\max}:=\max\{N_{i},i\in\mathcal{N}\}$.
@@
According to the definitions of $\mathbf{A}_{i}$ and $\mathbf{H}_{i}$ in equations (II-C) and (II-C), we have
@@
$\mathbf{A}_{i}^{T}\mathbf{A}_{i}=\begin{bmatrix}N_{i}&-\mathbf{1}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}&0&\mathbf{0}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}\\ -\mathbf{1}_{N_{i}}&\mathbf{I}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\\ \mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\\ 0&\mathbf{0}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}&2N_{i}&-\mathbf{1}_{N_{i}}^{T}&-\mathbf{1}_{N_{i}}^{T}\\ \mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}&-\mathbf{1}_{N_{i}}&\mathbf{I}_{N_{i}}&\mathbf{O}_{N_{i}}\\ \mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}&-\mathbf{1}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{I}_{N_{i}}\end{bmatrix}\otimes\mathbf{I}_{n}\in\mathbb{R}^{n\cdot(2+4N_{i})\times n\cdot(2+4N_{i})},$ (B.1)
@@
$\mathbf{H}_{i}^{T}\mathbf{H}_{i}=\begin{bmatrix}N_{i}+1&\mathbf{0}_{N_{i}}^{T}&-\mathbf{1}_{N_{i}}^{T}&-1&\mathbf{0}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}\\ \mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\\ -\mathbf{1}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{I}_{N_{i}}&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\\ -1&\mathbf{0}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}&1&\mathbf{0}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}\\ \mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\\ \mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\end{bmatrix}\otimes\mathbf{I}_{n}\in\mathbb{R}^{n\cdot(2+4N_{i})\times n\cdot(2+4N_{i})}.$ (B.2)
Substituting equations (B.1) and (B.2) into equation (32) yields
@@
$c\mathbf{B}_{i}^{T}\mathbf{B}_{i}=\begin{bmatrix}(c+1)N_{i}+1&c\cdot\mathbf{1}_{N_{i}}^{T}&\mathbf{1}_{N_{i}}^{T}&1&\mathbf{0}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}\\ c\cdot\mathbf{1}_{N_{i}}&c\cdot\mathbf{I}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\\ \mathbf{1}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{I}_{N_{i}}&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\\ 1&\mathbf{0}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}&2cN_{i}+1&c\cdot\mathbf{1}_{N_{i}}^{T}&c\cdot\mathbf{1}_{N_{i}}^{T}\\ \mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}&c\cdot\mathbf{1}_{N_{i}}&c\cdot\mathbf{I}_{N_{i}}&\mathbf{O}_{N_{i}}\\ \mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}&c\cdot\mathbf{1}_{N_{i}}&\mathbf{O}_{N_{i}}&c\cdot\mathbf{I}_{N_{i}}\end{bmatrix}\otimes\mathbf{I}_{n}\in\mathbb{R}^{n\cdot(2+4N_{i})\times n\cdot(2+4N_{i})}.$ (B.3)
Subtracting $\mathbf{z}_{i}=\mathbf{z}_{i}^{t}$ from $\mathbf{z}_{i}=\mathbf{z}_{i}^{t+1}$ in equation (III), we obtain
@@
Substituting equation (B.5) into equation (B), we derive
@@
According to the definition of $\mathbf{A}_{i}^{T}\mathbf{A}_{i}$ in equation (B.1), we have $\sigma_{\min}=1$. Combining this inequality with equation (B), we obtain
@@
$\leq$ $3(N_{i}+2)\|\mathbf{D}_{i}\left(\mathbf{w}_{i}^{t}-\mathbf{w}_{i}^{t-1}\right)-\mathbf{H}_{i}^{T}\left(\tilde{\mathbf{z}}_{i}^{t+1}-\tilde{\mathbf{z}}_{i}^{t}\right)\|^{2}$
$+3c^{2}(4N_{i}+2)\|\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{z}_{i}^{t+1}-(\tilde{\mathbf{z}}_{i}^{t}-\mathbf{z}_{i}^{t})\|_{\mathbf{A}_{i}^{T}\mathbf{A}_{i}}$
@@
The last inequality above holds based on equations (B.1)–(B.3), the triangle inequality of norms, and the facts $\|\mathbf{H}_{i}^{T}\mathbf{H}_{i}\|=N_{i}+2$, $\|\mathbf{A}_{i}^{T}\mathbf{A}_{i}\|=2N_{i}+1$. Here, $\|\mathbf{A}_{i}^{T}\mathbf{A}_{i}\|$ denotes the spectral norm of the matrix $\mathbf{A}_{i}^{T}\mathbf{A}_{i}$. Combining with the form of $c\mathbf{B}_{i}^{T}\mathbf{B}_{i}$ in equation (B.3), we have
$\|\mathbf{B}_{i}^{T}\mathbf{B}_{i}\|\leq\frac{\left(1+2c\right)\left(2+N_{\max}\right)}{c},$ (B.12)
where $N_{\max}=\max\left\{N_{i},i\in\mathcal{N}\right\}$. Substituting equation (B.12) into equation (B) and applying the result to equation (B.8), we get
@@
$\leq$ $\frac{3(N_{max}+2)}{c}\|\mathbf{D}_{i}\left(\mathbf{w}_{i}^{t}-\mathbf{w}_{i}^{t-1}\right)-\mathbf{H}_{i}^{T}\left(\tilde{\mathbf{z}}_{i}^{t+1}-\tilde{\mathbf{z}}_{i}^{t}\right)\|^{2}$
$+3c(2N_{max}+1)\|\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{z}_{i}^{t+1}-(\tilde{\mathbf{z}}_{i}^{t}-\mathbf{z}_{i}^{t})\|_{\mathbf{A}_{i}^{T}\mathbf{A}_{i}}$
$+3(1+2c)(2+N_{max})\|\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{z}_{i}^{t}-(\tilde{\mathbf{z}}_{i}^{t}-\mathbf{z}_{i}^{t-1})\|_{\mathbf{B}_{i}^{T}\mathbf{B}_{i}}$ (B.13)
Combining equations (B.6)–(B.7) and equation (B), and summing over nodes $i\in\mathcal{N}$, we derive inequality (4). ∎
@@
where $d_{\max}=\max\{r_{i},d_{i,j},i\in\mathcal{N},j\in\mathcal{N}_{i}\}$.
@@
From equation (B), we have:
@@
Next, we transform both sides of equation (B). First, the left-hand side of equation (B) can be expressed as:
@@
The last equality holds due to the identity: for any $a,b\in\mathbb{R}^{(4N_{i}+2)n}$,
$\langle a,b\rangle=-\frac{1}{2}\|a-b\|^{2}+\frac{1}{2}\|a\|^{2}+\frac{1}{2}\|b\|^{2}.$
@@
where $\|\mathbf{A}_{i}^{T}\mathbf{A}_{i}\|=2N_{i}+1$ is derived from equation (B.1).
Referring to Section 3 of [25], the right-hand side of equation (B) can be expressed as:
@@
Substituting equations (B.16) and (B.18) into equations (B) and (B) respectively, and again using $\mathbf{U}_{i}=\mathbf{H}_{i}^{T}\mathbf{H}_{i}+c\mathbf{A}_{i}^{T}\mathbf{A}_{i}+c\mathbf{B}_{i}^{T}\mathbf{B}_{i}$ together with equation (B), we can derive the final result. ∎
@@
$-\frac{c\cdot\tilde{\tau}_{\min}}{6N_{sum}n(c+1)^{2}}\|\mathbf{z}_{i}^{t}-\tilde{\mathbf{z}}_{i}^{t+1}\|^{2}\Bigg],$
@@
$\tilde{\tau}_{min}:=\min\quad\{5(c+1)^{2}N_{i}^{2}+(3c^{2}+4c+3)N_{i},i\in\mathcal{N}\},\quad N_{\text{sum}}:=\sum_{i\in\mathcal{N}}N_{i}.$
@@
Using equation (B) again, we obtain
@@
First, the left-hand side of equation (B) can be written as:
@@
The first equality holds due to equation (26), and the second equality follows from the fact that for any $\mathbf{a},\mathbf{b}\in\mathbb{R}^{(4N_{i}+2)n}$, we have
@@
For the right-hand side of equation (B), using equation (B), we obtain

### v2

@@ 6 Conclusion
{ack}
Partially supported by the Roman Senate.
A.1 Solution to Remark 2
@@
By substituting the explicit form of equation (3) into problem (30), the optimization problem (30) can be reformulated as:
$\min_{\mathbf{z}\in\mathcal{X}}\sum_{i\in\mathcal{N}}\Big($ $[(c+1)N_{i}+M_{i}]\|\mathbf{x}_{i}-\tilde{\mathbf{x}}_{i}^{t+1}\|^{2}+\sum_{j\in\mathcal{N}_{i}}c\|\mathbf{p}_{i,j}^{-}-(\tilde{\mathbf{p}}_{i,j}^{-})^{t+1}\|^{2}$
$+\sum_{j\in\mathcal{N}_{i}}\|\mathbf{p}_{i,j}^{+}-(\tilde{\mathbf{p}}_{i,j}^{+})^{t+1}\|^{2}+\sum_{k\in\mathcal{M}}((2cN_{i}+\mathbb{I}_{\mathcal{M}_{i}}(k))\|\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}-\tilde{\mathbf{y}}_{i,{\color[rgb]{0,0,0}k}}^{t+1}\|^{2}$
$+\sum_{j\in\mathcal{N}_{i}}c\|\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{-}-(\tilde{\mathbf{q}}_{i,j,{\color[rgb]{0,0,0}k}}^{-})^{t+1}\|^{2}+\sum_{j\in\mathcal{N}_{i}}c\|\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{+}-(\tilde{\mathbf{q}}_{i,j,{\color[rgb]{0,0,0}k}}^{+})^{t+1}\|^{2})\Big).$
@@
The optimal solution for $\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t+1}$ is:
$\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t+1}=\tilde{\mathbf{y}}_{i,{\color[rgb]{0,0,0}k}}^{t+1},\text{for }i\in\mathcal{N}$
@@
The terms related to $\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{-}$ are:
$c\|\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{-}-(\tilde{\mathbf{q}}_{i,j,{\color[rgb]{0,0,0}k}}^{-})^{t+1}\|^{2}+c\|\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{-}-(\tilde{\mathbf{q}}_{j,i,{\color[rgb]{0,0,0}k}}^{+})^{t+1}\|^{2},\quad\forall i\in\mathcal{N},j\in\mathcal{N}_{i},k\in\mathcal{M}.$
The terms related to $\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{+}$ are:
$c\|\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{+}-(\tilde{\mathbf{q}}_{i,j,{\color[rgb]{0,0,0}k}}^{-})^{t+1}\|^{2}+c\|\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{+}-(\tilde{\mathbf{q}}_{j,i,{\color[rgb]{0,0,0}k}}^{+})^{t+1}\|^{2},\quad\forall i\in\mathcal{N},j\in\mathcal{N}_{i},k\in\mathcal{M}.$
@@
For $i\in\mathcal{N},j\in\mathcal{N},k\in\mathcal{M}$, the closed-form solution for $\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{+}$ is:
$(\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{-})^{t+1}=\frac{1}{2}\left((\tilde{\mathbf{q}}_{i,j,{\color[rgb]{0,0,0}k}}^{-})^{t+1}+(\tilde{\mathbf{q}}_{j,i,{\color[rgb]{0,0,0}k}}^{+})^{t+1}\right),\quad\forall i\in\mathcal{N},j\in\mathcal{N}_{i}.$
And the closed-form solution for $\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{-}$ is:
$(\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{+})^{t+1}=\frac{1}{2}\left((\tilde{\mathbf{q}}_{i,j,{\color[rgb]{0,0,0}k}}^{+})^{t+1}+(\tilde{\mathbf{q}}_{j,i,{\color[rgb]{0,0,0}k}}^{-})^{t+1}\right),\quad\forall i\in\mathcal{N},j\in\mathcal{N}_{i}.$
A.2 (Proof of Lemma1)
@@
$\quad+\frac{3\left(N_{\max}+M_{\max}+1\right)}{c}\|\mathbf{H}_{i}\Delta\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{D}_{i}\Delta\mathbf{w}_{i}^{t}\|^{2}$
$\quad+3c^{2}\left(2N_{\max}+1\right)\|\Delta\tilde{\mathbf{z}}_{i}^{t+1}-\Delta\mathbf{z}_{i}^{t+1}\|_{\mathbf{A}_{i}^{T}\mathbf{A}_{i}}^{2}$
$\quad+3c\left(1+c\right)\left(1+M_{\max}+2N_{\max}\right)\|\Delta\tilde{\mathbf{z}}_{i}^{t+1}-\Delta\mathbf{z}_{i}^{t}\|_{\mathbf{B}_{i}^{T}\mathbf{B}_{i}}^{2}\Bigg],$ (B.1)
where $N_{\max}:=\max\{N_{i},i\in\mathcal{N}\},M_{\max}:=\max\{M_{i},i\in\mathcal{M}\}$.
@@
According to the definitions of $\mathbf{A}_{i}$ and $\mathbf{H}_{i}$ in equations (19) and (17), we have
$\mathbf{A}_{i}^{T}\mathbf{A}_{i}$ $=\operatorname{blockdiag}\left({\mathbf{A}_{i}^{\text{sen}}}^{T}\mathbf{A}_{i}^{\text{sen}},{\mathbf{A}_{i,1}^{\text{tar}}}^{T}\mathbf{A}_{i,1}^{\text{tar}},\dots,{\mathbf{A}_{i,M}^{\text{tar}}}^{T}\mathbf{A}_{i,M}^{\text{tar}}\right)\otimes\mathbf{I}_{n}$
$\in\mathbb{R}^{n\cdot(2+4N_{i})\times n\cdot(2+4N_{i})},$ (B.2)
where,
@@
${\mathbf{A}_{i}^{\text{sen}}}^{T}\mathbf{A}_{i}^{\text{sen}}=\begin{bmatrix}N_{i}&-\mathbf{1}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}\\ -\mathbf{1}_{N_{i}}&\mathbf{I}_{N_{i}}&\mathbf{O}_{N_{i}}\\ \mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\end{bmatrix}\otimes\mathbf{I}_{n}\in\mathbb{R}^{n\cdot(2+4N_{i})\times n\cdot(2+4N_{i})},$
and $\forall k\in\mathcal{M}$,
@@
${\mathbf{A}_{i,{\color[rgb]{0,0,0}k}}^{\text{tar}}}^{T}\mathbf{A}_{i,{\color[rgb]{0,0,0}k}}^{\text{tar}}=\begin{bmatrix}2N_{i}&-\mathbf{1}_{N_{i}}^{T}&-\mathbf{1}_{N_{i}}^{T}\\ -\mathbf{1}_{N_{i}}&\mathbf{I}_{N_{i}}&\mathbf{O}_{N_{i}}\\ -\mathbf{1}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{I}_{N_{i}}\end{bmatrix}\otimes\mathbf{I}_{n}\in\mathbb{R}^{n\cdot(2+4N_{i})\times n\cdot(2+4N_{i})};$
@@
$\mathbf{H}_{i}^{T}\mathbf{H}_{i}$ $=\begin{bmatrix}N_{i}+M_{i}&\mathbf{0}_{N_{i}}^{T}&-\mathbf{1}_{N_{i}}^{T}&-{\color[rgb]{0,0,0}\mathbb{I}_{\mathcal{M}_{i}}(1)}&\mathbf{0}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}&...&{\color[rgb]{0,0,0}\mathbb{I}_{\mathcal{M}_{i}}(M)}&\mathbf{0}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}\\ \mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}&...&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\\ -\mathbf{1}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{I}_{N_{i}}&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}&...&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\\ -{\color[rgb]{0,0,0}\mathbb{I}_{\mathcal{M}_{i}}(1)}&\mathbf{0}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}&{\color[rgb]{0,0,0}\mathbb{I}_{\mathcal{M}_{i}}(1)}&\mathbf{0}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}&...&0&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\\ \mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}&...&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\\ \mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}&...&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ -{\color[rgb]{0,0,0}\mathbb{I}_{\mathcal{M}_{i}}(M)}&\mathbf{0}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}&0&\mathbf{0}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}&...&{\color[rgb]{0,0,0}\mathbb{I}_{\mathcal{M}_{i}}(M)}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\\ \mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}&...&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\\ \mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}&...&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\end{bmatrix}\otimes\mathbf{I}_{n}$
$\in\mathbb{R}^{n\cdot(2+4N_{i})\times n\cdot(2+4N_{i})}.$ (B.3)
Subtracting $\mathbf{z}_{i}=\mathbf{z}_{i}^{t}$ from $\mathbf{z}_{i}=\mathbf{z}_{i}^{t+1}$ in equation (3), we obtain
@@
Substituting equation (B.5) into equation (A.2), we derive
@@
According to the definition of $\mathbf{A}_{i}^{T}\mathbf{A}_{i}$ in equation (A.2), we have $\sigma_{\min}=1$. Combining this inequality with equation (A.2), we obtain
@@
$\leq$ $3(N_{i}+M_{i}+1)\|\mathbf{D}_{i}\left(\mathbf{w}_{i}^{t}-\mathbf{w}_{i}^{t-1}\right)-\mathbf{H}_{i}^{T}\left(\tilde{\mathbf{z}}_{i}^{t+1}-\tilde{\mathbf{z}}_{i}^{t}\right)\|^{2}$
$+3c^{2}(2N_{i}+1)\|\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{z}_{i}^{t+1}-(\tilde{\mathbf{z}}_{i}^{t}-\mathbf{z}_{i}^{t})\|_{\mathbf{A}_{i}^{T}\mathbf{A}_{i}}$
@@
The last inequality above holds based on equations (A.2),(A.2), the triangle inequality of norms, and the facts $\|\mathbf{H}_{i}^{T}\mathbf{H}_{i}\|=N_{i}+M_{i}+1$, $\|\mathbf{A}_{i}^{T}\mathbf{A}_{i}\|=2N_{i}+1$. Here, $\|\mathbf{A}_{i}^{T}\mathbf{A}_{i}\|$ denotes the spectral norm of the matrix $\mathbf{A}_{i}^{T}\mathbf{A}_{i}$. Combining with the form of $c\mathbf{B}_{i}^{T}\mathbf{B}_{i}$ , we have
$\|\mathbf{B}_{i}^{T}\mathbf{B}_{i}\|\leq\frac{\left(1+c\right)\left(1+M_{\max}+2N_{\max}\right)}{c},$ (B.12)
where $N_{\max}=\max\left\{N_{i},i\in\mathcal{N}\right\},M_{\max}=\max\left\{M_{i},i\in\mathcal{M}\right\}$. Substituting equation (B.12) into equation (A.2) and applying the result to equation (B.8), we get
@@
$\leq$ $\frac{3(N_{max}+M_{max}+2)}{c}\|\mathbf{D}_{i}\left(\mathbf{w}_{i}^{t}-\mathbf{w}_{i}^{t-1}\right)-\mathbf{H}_{i}^{T}\left(\tilde{\mathbf{z}}_{i}^{t+1}-\tilde{\mathbf{z}}_{i}^{t}\right)\|^{2}$
$+3c^{2}(2N_{max}+1)\|\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{z}_{i}^{t+1}-(\tilde{\mathbf{z}}_{i}^{t}-\mathbf{z}_{i}^{t})\|_{\mathbf{A}_{i}^{T}\mathbf{A}_{i}}$
$+3(1+c)\left(1+M_{\max}+2N_{\max}\right)\|\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{z}_{i}^{t}-(\tilde{\mathbf{z}}_{i}^{t}-\mathbf{z}_{i}^{t-1})\|_{\mathbf{B}_{i}^{T}\mathbf{B}_{i}}$ (B.13)
Combining equations (B.6)–(B.7) and equation (A.2), and summing over nodes $i\in\mathcal{N}$, we derive inequality (4). ∎
@@
where $d_{\max}=\max\left(\{d_{i,j}\mid i\in\mathcal{N},j\in\mathcal{N}_{i}\}\cup\{r_{i,{\color[rgb]{0,0,0}k}}\mid i\in\mathcal{N},k\in\mathcal{M}_{i}\}\right)$.
@@
From equation (A.2), we have:
@@
Next, we transform both sides of equation (A.2). First, the left-hand side of equation (A.2) can be expressed as:
@@
The last equality holds due to the identity: for any $\mathbf{a},\mathbf{b}\in\mathbb{R}^{(2N_{i}+1)(M+1)n}$,
$\langle\mathbf{a},\mathbf{b}\rangle=-\frac{1}{2}\|\mathbf{a}-\mathbf{b}\|^{2}+\frac{1}{2}\|\mathbf{a}\|^{2}+\frac{1}{2}\|\mathbf{b}\|^{2}.$
@@
where $\|\mathbf{A}_{i}^{T}\mathbf{A}_{i}\|=2N_{i}+1$ is derived from equation (A.2).
The right-hand side of equation (A.2) can be expressed as:
@@
Substituting equations (B.16) and (B.18) into equations (A.2) and (A.2) respectively, and again using $\mathbf{U}_{i}=\mathbf{H}_{i}^{T}\mathbf{H}_{i}+c\mathbf{A}_{i}^{T}\mathbf{A}_{i}+c\mathbf{B}_{i}^{T}\mathbf{B}_{i}$ together with equation (A.2), we can derive the final result. ∎
@@
$-\frac{c\cdot\tilde{\tau}_{\min}}{2(2M+1)N_{\text{sum}}n(c+1)^{2}}\|\mathbf{z}_{i}^{t}-\tilde{\mathbf{z}}_{i}^{t+1}\|^{2}\Bigg],$
@@
$\tilde{\tau}_{min}:=\min\quad\{(c+1)^{2}[(6M+1)N_{i}^{2}+N_{i}],i\in\mathcal{N}\},\quad N_{\text{sum}}:=\sum_{i\in\mathcal{N}}N_{i}.$
@@
Using equation (A.2) again, we obtain
@@
First, the left-hand side of equation (A.2) can be written as:
@@
The first equality holds due to equation (25), and the second equality follows from the fact that for any $\mathbf{a},\mathbf{b}\in\mathbb{R}^{(2N_{i}+1)(M+1)n}$, we have
@@
For the right-hand side of equation (A.2), using equation (A.2), we obtain